ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

3. Number Theory and Arithmetic Algebraic Geometry

1991 MS Classification: 11D41, 11D68 Tirado, Eleazar Towards an Elementary Approach to Fermats Last Theorem Work represents effort done since 1987. Effort was conducted with the only purpose of finding a true elementary way to deal with FsLT. Keys: Fermats Little Theorem, uniqueness of Newtons polynomial expression, theorems 48 and 72(generalized): Shanks. D. “Solved and Unsolved Problems in Number Theory. ” Chelsea, NY.1978. Parametric equations have been found through which all possibili ties for odd terms, (x, y), of equation are covered and the even be comes Z=f(x, y), with built in possibilities: (xy)=even pth power First Case[p(prime exponent) doesnt divide any term or p divides one of the odd] and p(xy)=even pth power-Second Case[p divides even term: xy)]. These set Conjecture in worst scenary to be true.